# MC-NeRF: Multi-Camera Neural Radiance Fields for Multi-Camera Image Acquisition Systems

### v2

@@ 4.1   Dataset and Hardware Platfrom
In this experiment, our goal is to verify the effectiveness of the proposed intrinsic parameters regression method. We conduct experiments using AprilTag images collected by 70 cameras with different FOVs. The dataset consists of three groups: the images in the first group contain only one AprilTag in each image, denoted as $Dataset\_A$. This group ensures that the calibration points lie in the same plane. The images in the second group contain two AprilTags, denoted as $Dataset\_B$. According to the conclusion in Section 3.2, theoretically, the results trained by this dataset can successfully obtain the intrinsic parameters. The images in the third group contain three AprilTags, denoted as $Dataset\_C$. We believe that the results trained with this group will have better accuracy. Here, $Fx$ and $Fy$ are the focal lengths of the camera in the horizontal and vertical directions, while $ux$ and $uy$ are the coordinates of the principal point.
@@
The regression results are presented in Fig. 9 and Table 3. Average loss of intrinsic parameters refers to the average absolute error between predicted intrinsic parameters and ground truth intrinsic parameters. In this scenario, the images contain only 1 AprilTag ($Dataset\_A$) and the intrinsic parameters cannot be obtained. This case exhibits similar performance to the degenerate case, indicating the persistence of degeneracy in the proposed method. Specifically, as illustrated by the black curves in $E2$ and $E3$, $Fx$ and $Fy$ fail to converge, and the losses in $E4$ and $E5$ exhibit significant deviations from 00. Moreover, Table 3 reveals that the accuracy of the regression results is significantly inferior to the other two cases.
@@
When the images we used contain more than two AprilTags ($Dataset\_B$, $Dataset\_C$), the regression can be guaranteed to converge. It can be observed from the quantified data in the table that the parameters trained using the dataset containing three AprilTags achieve higher accuracy, as demonstrated in $Fx$, $Fy$, $ux$, and $uy$. In comparison, the parameters trained with the dataset containing 2 AprilTags converge more rapidly.
In summary, camera intrinsic parameters can be obtained when the Intrinsic Reprojection Loss branch is trained with dataset containing at least two Apriltags. This demonstrates the effectiveness of the proposed method. If the provided calibration points lie on the same plane, the degenerated case will still occur, resulting in the inability to obtain camera intrinsic parameters. The degenerated case is reproducible in our method.
4.4   Exp.3 Extrinsic Parameters Regression
@@
Figure 10: Comparison of poses regression performance be- tween BARF and L2G-NeRF with the random initialization poses.
TABLE 4: Comparison of poses regression performance between BARF and L2G-NeRF with the initialization poses obtained from calibration cube.
@@
In this experiment, we aim to verify the necessity of extrinsic parameters regression. While BARF and L2G-NeRF address the joint optimization of NeRF and camera poses, these methods are limited by the initial extrinsic parameters. Specifically, both of the above methods add perturbations to the ground truth extrinsic parameters, leading to camera poses drift, which is then attempted to be rectified during the joint optimization. However, the initial pose of each camera in practical applications still requires huge preparation work, especially for the multi-camera acquisition systems with a large number of cameras. Therefore, in this experiment we randomly provide the initial poses of all cameras and attempt to perform poses regression on four styles of scenes using the above mentioned two methods. This is done to assess the performance of the previous methods under the condition where no potential poses information is provided.
Figure 11: Qualitative results of the planar image alignment experiment. We visualize the optimized warps (top row), the patch reconstructions in corresponding colors (middle row), and recovered image representation from ${f_{mlp}}$ (bottom row). MC-NeRF with calibration points is able to recover accurate alignment and high-fidelity image reconstruction
The results are presented in Fig. 10. The first row illustrates the randomly initialized camera poses of the two methods across the four scenarios, while the second row depicts the stabilized camera poses after training. When the camera poses are initialized randomly, both methods struggle to acquire accurate camera extrinsic parameters. As our proposed method employs the progressive alignment technique introduced by BARF in the joint optimization, it becomes evident that an effective initialization of the camera pose is essential. Further details regarding the convergence process are provided in our supplementary project webpage.
It is worth noting that the camera initialization pose obtained from the calibration cube can satisfy both L2G-NeRF and BARF. The reasons for adopting the BARF approach are as follows: Firstly, as outlined in the original BARF, it makes only minor modifications to the classic Coarse-to-Fine architecture of NeRF, which makes it easily extensible. Secondly, in comparison to L2G-NeRF, these two methods exhibit similar performance under the aforementioned extrinsic initialization conditions. The comparisons are presented in Table 4. Inherited Value refers to the initial poses obtained by calibration cube, $Loss\_R$ and $Loss\_T$ denote the losses associated with the ground truth rotation and translation data. These losses are defined as follows:
@@
${\hat{R}_{i}}$ and ${R_{i}}$ represent the ground truth and predicted rotation matrices respectively, while ${\hat{T}_{i}}$ and ${T_{i}}$ represent the ground truth and predicted translation matrices. $mean()$ represents the mean function. $n$ represents the total number of poses, which is equal to the number of images in the dataset. It can be observed that when using the initial pose provided by the calibration cube, there is almost no difference in performance between these methods across the four different styles of scenes.
@@
In summary, extrinsic parameter regression is necessary. Randomly initializing external parameters cannot guarantee obtaining accurate results. By using calibration cube data to initialize extrinsic parameters, we can subsequently achieve precise camera poses using the previously mentioned methods.
@@
[…]
BARF 85.6157 14.2270 0.5149 0.6641
L2G-NeRF 55.3545 20.5440 0.7884 0.3364
MC-NeRF(w/o) 55.3096 15.9662 0.5923 0.6268
MC-NeRF(ours) 0.4084 40.8309 0.9828 0.0318
4.5   Exp.4 Planar Image Alignment (2D)
In this experiment, our goal is to compare the performance of the proposed method with previous works in the task of 2D image alignment. Furthermore, we also explore the performance of the method when calibration points are not provided. We choose the same representative image "cat" as BARF used from ImageNet[51]. We select five patches sampled from the original image, apply linear transformations and rigid perturbations, and optimize Eq.(17) to find the transformation $L$ and $T$ for each patch. Simultaneously, we train the neural network ${f_{mlp}}$ to represent the entire image.
@@
The initial poses and ground truth transformations for each patch are depicted on the left side of Fig. 11. It’s important to note that in BARF, MC-NeRF (W/O), and MC-NeRF, $L$ represents linear transformation, and $T$ represents rigid rotation and translation. However, in L2G-NeRF, we combine $L$ and $T$ and directly replace them with a homography transformation. The reason for this lies in the design of L2G-NeRF, where finding differentiable parameter estimation solvers for custom computational flows is challenging. In theory, the homography transformation solver provided by L2G-NeRF itself can simultaneously handle scale, rotation, and translation, and we use the homography transformation to replace $L$ and $T$.
As mentioned in Section 3, joint optimization of intrinsic parameters can be equivalent to adding an additional linear transformation $L$. To decouple this linear transformation from rotation and translation (as discussed in Section 3.4.2), we selected six pairs of randomly chosen calibration points as constraints for each patch. This constraint is analogous to the calibration points obtained by calibration cube in 3D dimensional space. In 2D space, the loss function for MC-NeRF is $Eq.(20)$. In all test models, a 6-layer fully connected network with 256 neurons each is respectively defined. The training epochs are set to 50,000 to ensure convergence for all methods. The specific regression process is demonstrated on our project webpage.
TABLE 6: Quantitative rendering results of Fix-Step NeRF and MC-NeRF on all scenes.
TABLE 7: Quantitative camera parameters regression results of Fix-Step NeRF and MC-NeRF on all scenes.

### v4

@@ 4   EXPERIMENTS
In this experiment, we aim to verify whether the proposed auxiliary image capture scheme can avoid the degeneration case. We conduct experiments using AprilTag images collected by 70 cameras with different FOVs. The dataset consists of three groups: the images in the first group contain only one AprilTag in each image, denoted as $Dataset\_A$. This group ensures that the selected 3D points lie in the same plane. The images in the second group contain two AprilTags, denoted as $Dataset\_B$. According to the conclusion in Sec 3.3.1, the results trained by this dataset can successfully obtain the intrinsics and avoid degeneration case. The images in the third group contain three AprilTags, denoted as $Dataset\_C$. We believe that the results trained with this group will have better accuracy. Here, $Fx$ and $Fy$ are the focal lengths of the camera in the horizontal and vertical directions, while $ux$ and $uy$ are the coordinates of the principal point.
@@
The results are presented in Fig. 10 and Table 2. In Fig. 10, average loss of intrinsics refers to the average absolute error between predicted intrinsics and ground truth. In this scenario, the intrinsics cannot be obtained with the images contain only 1 AprilTag ($Dataset\_A$). This case exhibits performance to the degeneration case. Specifically, as illustrated by the black curves in $E2$ and $E3$, $Fx$ and $Fy$ fail to converge, and the losses in $E4$ and $E5$ exhibit significant deviations from 00. Moreover, Table 2 reveals that the accuracy of the results is significantly inferior to the other two cases.
@@
When the images we used contain more than two AprilTags ($Dataset\_B$, $Dataset\_C$), the estimation can be guaranteed to converge. It can be observed from the quantified data in the table that the parameters trained using the dataset containing three AprilTags achieve higher accuracy, as demonstrated in $Fx$, $Fy$, $ux$, and $uy$. In comparison, the parameters trained with the dataset containing 2 AprilTags converge more rapidly.
In summary, the image capture scheme we proposed is effective and can prevent the degeneration case. If the provided 3D points lie on the same plane, the degeneration case will occur. Including more than two AprilTags in a single auxiliary image can provide effective intrinsic constraints, ensuring that the joint optimization process yields accurate results.
4.3   Exp.3 Initialization of Extrinsics Estimation
1
@@
Figure 11: Comparison of extrinsics estimation performance between BARF and L2G-NeRF with the random initialization values.
In this experiment, we aim to verify the necessity of extrinsics estimation with $Pack1$. Both BARF and L2G-NeRF address the joint optimization of NeRF and extrinsics by adding perturbations to the ground truth extrinsics. This leads to camera pose drift, which the methods then attempt to rectify during the joint optimization process.
However, the initial extrinsics of each camera in practical applications still requires huge preparation work, especially for the multi-camera acquisition systems with a large number of cameras. Therefore, in this experiment we randomly provide the initial extrinsics of all cameras and attempt to perform extrinsics estimation on four styles of scenes using the above mentioned two methods. This is done to assess the performance of the previous methods under the condition where no potential poses information is provided.
The results are presented in Fig. 11. The first row illustrates the randomly initialized extrinsics of the two methods across the four scenarios, while the second row depicts the stabilized extrinsics after training. When the extrinsics are initialized randomly, both methods struggle to acquire accurate values. As our proposed method employs the progressive alignment technique introduced by BARF in the joint optimization, it becomes evident that an effective initialization of the extrinsics is essential. Further details regarding the convergence process are provided in our supplementary project webpage.
TABLE 3: Comparison of extrinsics estimation performance between BARF and L2G-NeRF with the initialization extrinsics obtained from AprilTag cube. $Inherited$ $Value$ represents the initial camera position obtained directly using the reprojection error method.
@@
Figure 12: Rendering results of different methods in real-world dataset. We select the $Lab1$ scene from proposed real-world dataset to showcase rendering quality, while the results for the remaining two scenes are provided in the supplementary materials and on project website. In our comparative experiments, InstantSplat and CF Gaussian autonomously estimate camera parameters without any provided initializations. Methods such as CAMP, NeRFacto, and BARF, which are initialized using camera parameters provided by Colmap, are limited to optimizing only the extrinsic parameters. In contrast, SC-NeRF and NeRF--- -, also initialized with Colmap, are capable of optimizing both intrinsic and extrinsic parameters. Instant-NGP, which lacks camera parameter optimization capabilities, serves as a baseline reference. MC-NeRF utilizes auxiliary images for initialization, and it is also capable of optimizing both intrinsic and extrinsic parameters. The results demonstrate that our method achieved the competitive rendering performance among all compared methods.
Random initialization of extrinsics can lead to optimization failure. We provide auxiliary images from $Pack1$ to assist in obtaining initial values of the extrinsics. With these initial values provided, the performance of the above two methods is shown in Table 3. Inherited Value refers to the initial values provided by AprilTag cube, $Loss\_R$ and $Loss\_T$ denote the losses associated with the ground truth rotation and translation data. These losses are defined as follows:
@@
${\hat{R}_{i}}$ and ${R_{i}}$ represent the ground truth and predicted rotation matrices respectively, while ${\hat{T}_{i}}$ and ${T_{i}}$ represent the ground truth and predicted translation matrices. $mean()$ represents the mean function. $n$ represents the total number of poses, which is equal to the number of images in the dataset. It can be observed that when using the initial intrinsics, these methods achieved high performance under all four styles.
@@
In summary, it is necessary to use the auxiliary image set $Pack1$ to provide initial extrinsics for the camera. Randomly initializing extrinsics cannot guarantee obtaining accurate results. By using AprilTag cube to obtain initial values, we can subsequently achieve precise extrinsics results.
4.4   Exp.4 Comparison with Other methods
In this experiment, we aim to validate the effectiveness of our method in real-world scenarios. We selecte NeRF series methods, including CAMP, NeRF--- -, BARF, SC-NeRF, Nerfacto and Instant-NGP as baselines. Additionally, we included 3D Gaussian-based methods, specifically CF-3D Gaussian and InstantSplat, as additional baselines. For methods that require initial camera parameters, we rely on the widely used COLMAP for initialization. In contrast, our approach incorporates auxiliary images to assist in this process. The real-world dataset consists of three scenes: $Lab1$, $Lab2$, and $Lab3$. Due to space limitations, we present the training results of $Lab1$ in the main manuscript, while the rendered results for the other two scenes are provided in the supplementary materials.
The experimental results are presented in Fig.12, Table 4, and Table 5. Regarding image rendering quality, in Fig.12, we select six views generated by different methods randomly, and the results demonstrate that our method achieves competitive performance. In this experiment, INGP is used as the baseline without any camera parameter optimization. Only SC-NeRF, NeRF--- -, and our method jointly optimize both the intrinsic and extrinsic parameters, while the remaining methods optimize only the extrinsic parameters. It is important to note that NeRF--- - is the only method that failed to render in all cases. In the original paper, the authors explicitly state that their approach is only applicable to forward-facing scenes, which is consistent with our experimental results.
TABLE 4: Quantitative camera parameters estimation results of different methods on real-world datasets. The table shows that our method achieves competitive results in all three aspects: camera intrinsics, camera extrinsics, and scene reconstruction scale.
Scene Method $Loss\_Fx\downarrow$ $Loss\_Fy\downarrow$ $Loss\_Cx\downarrow$ $Loss\_Cx\downarrow$ $Loss\_R\downarrow$ $Loss\_T\downarrow$ $Loss\_Scale\uparrow$
Lab 1 InstantSplat 20.2304 20.2586 8.3652 8.7434 0.2619 0.0982 0.0478
CF Gaussian 79.1077 79.1313 7.5821 8.4547 0.4124 1.8820 0.7239
CAMP 15.7198 15.5267 7.3414 8.5711 0.0258 0.0182 0.2195
NeRFacto 7.3588 7.1762 11.3652 11.7434 0.2679 0.0588 0.7946
BARF 7.3588 7.1762 11.3652 11.7434 0.2661 0.8986 0.2596
SC-NeRF 9.3722 13.0315 4.8873 9.3619 0.0135 0.0373 0.7991
NeRF--- - 163.1592 165.1681 18.3652 18.7434 0.4554 0.0177 0.0839
NeRF 7.3588 7.1762 11.3652 11.7434 0.2721 0.0749 0.3839
Instant-NGP 7.3588 7.1762 11.3652 11.7434 0.2721 0.0749 0.3839
MC-NeRF (ours) 4.1757 4.1002 3.8026 4.2973 0.0011 0.0129 0.9625
Lab 2 InstantSplat 19.4612 19.4894 8.3652 8.7434 0.2524 0.2459 0.0504
CF Gaussian 79.1077 79.1313 7.5821 8.4547 0.4328 2.0865 0.7155
CAMP 16.8892 16.8645 7.3414 8.5711 0.0089 0.2135 0.2224
NeRFacto 8.9282 8.4323 12.3652 12.7434 0.2381 0.2415 0.8015
BARF 8.9282 8.4323 12.3652 12.7434 0.2384 1.0924 0.2912
SC-NeRF 8.5852 15.4368 4.9523 9.1950 0.0165 0.2212 0.7947
NeRF--- - 176.5705 160.6003 18.3652 18.7434 0.4405 0.1968 0.0805
NeRF 8.9282 8.4323 12.3652 12.7434 0.2424 0.5393 0.3902
Instant-NGP 8.9282 8.4323 12.3652 12.7434 0.2424 0.5393 0.3902
MC-NeRF (ours) 3.9113 3.8856 4.1121 4.0004 0.0067 0.0286 0.9754
Lab 3 InstantSplat 19.2140 19.2423 8.3652 8.7434 0.2881 0.0557 0.0477
CF Gaussian 79.1077 79.1313 7.5821 8.4547 0.4016 1.9288 0.7345
CAMP 19.2211 19.3228 7.3414 8.5711 0.0028 0.0435 0.2173
NeRFacto 7.3223 7.3193 12.8605 12.5431 0.2466 0.0720 0.7861
BARF 7.3223 7.3193 12.8605 12.5431 0.2453 0.8915 0.2552
SC-NeRF 11.9082 15.5652 4.5782 9.3252 0.0159 0.0561 0.7914
NeRF--- - 129.3443 110.5904 10.6957 12.1557 0.4553 0.0102 0.1024
NeRF 7.3223 7.3193 12.8605 12.5431 0.2495 0.1509 0.3813
Instant-NGP 7.3223 7.3193 12.8605 12.5431 0.2495 0.1509 0.3813
MC-NeRF (ours) 4.2264 4.1906 4.9588 4.0681 0.0015 0.0341 0.9755
Table 4 presents the errors between the estimated camera parameters and the ground truth after joint optimization. Specifically, $loss\_Fx$ and $loss\_Fy$ represent the focal length errors, while $Loss\_Cx$ and $Loss\_Cy$ denote the principal point errors. $Loss\_R$ and $Loss\_T$ correspond to the rotational and translational errors of the extrinsic camera parameters. $Loss\_Scale$ represents the scale loss between the reconstructed scene and the ground truth scene, which is defined as follows:
$Loss\_Scale=\frac{{max(dist(T_{p}^{i}-T_{p}^{j}))}}{{max(dist(T_{g}^{i}-T_{g}^{j}))}}$ (21)
Where $T_{g}$ represents the translation vector of the ground truth camera pose, and $T_{p}$ denotes the translation vector of the estimated camera pose. $dist$ indicates the Euclidean distance, and $i$,$j$ representing the camera indices. $Loss\_Scale$ reflects the scale ratio between the reconstructed scene and the real-world scene. Unlike other methods that rely on equivalent camera models, MC-NeRF leverages auxiliary images to incorporate real-world scale, enabling the reconstruction of a 1:1 scale model of the scene. Additionally, our method achieves more accurate intrinsic camera parameters by using constraints from the auxiliary images, effectively constraining the solution space and allowing for the joint optimization of each camera’s corresponding parameters. As shown in the table, our approach outperforms others in terms of intrinsics and extrinsics estimation, as well as scale accuracy.
Table 5 shows the rendering performance metrics for each scene. Each scene is equipped with 88 cameras, from which we randomly selected 80 for the training dataset, while the remaining 8 were used for the test dataset. The metrics provided are calculated based on the rendering outcomes from these 8 test datasets. As demonstrated in the table, our method achieves the best performance in terms of rendering quality. Compared to other joint optimization methods, our work primarily focuses on reducing the barriers to implementing 3D reconstruction in multi-camera systems, rather than enhancing rendering quality. In MC-NeRF, the rendering model remains the classic NeRF framework. We believe the performance of our method benefits more from accurate camera poses than from the NeRF technique itself, which differs significantly from methods such as CAMP and the 3DGS series methods.
TABLE 5: Quantitative rendering results of different methods on real-world datasets. Compared to other methods, our approach also delivers competitive results in rendering quality. Unlike other works, our method retains the classic NeRF architecture, and we believe the improvement in rendering quality is driven by more accurate camera parameters rather than modifications to the NeRF model itself.
Scene Lab 1 Lab 2 Lab 3
$PSNR\uparrow$ $SSIM\uparrow$ $LPIPS\downarrow$ $PSNR\uparrow$ $SSIM\uparrow$ $LPIPS\downarrow$ $PSNR\uparrow$ $SSIM\uparrow$ $LPIPS\downarrow$
InstantSplat 20.3433 0.6086 0.3832 22.9553 0.7397 0.4981 20.5839 0.6233 0.4546
CF Gaussian 15.7241 0.4878 0.5872 16.5323 0.5314 0.5491 15.4723 0.4761 0.6001
CAMP 19.4850 0.5926 0.3714 20.7756 0.6816 0.3507 19.6813 0.5967 0.3654
NeRFacto 21.3255 0.6801 0.3963 21.6011 0.7304 0.3517 22.3201 0.7002 0.3489
BARF 18.4214 0.5645 0.7460 20.0294 0.6585 0.6673 20.8126 0.6342 0.5901
SC-NeRF 22.3175 0.6555 0.5272 23.8036 0.7369 0.4795 22.6125 0.6649 0.5288
NeRF--- - 13.4507 0.3778 0.8845 10.8519 0.4552 0.8529 14.0315 0.4027 0.8561
NeRF 22.3175 0.6555 0.5272 23.8036 0.7369 0.4795 24.0208 0.6935 0.4542
Instant-NGP 21.0498 0.6097 0.4174 23.1911 0.7202 0.3428 20.9077 0.6087 0.4113
MC-NeRF(ours) 23.1187 0.7199 0.3265 23.9901 0.7012 0.3277 24.3339 0.6923 0.3589
5   CONCLUSIONS
In this work, we present Multi-Camera Neural Radiance Fields (MC-NeRF), a 3D representation method designed for multi-camera image acquisition systems, which can joint optimize both camera parameters and NeRF. MC-NeRF breaks the assumption of a unique camera in previous studies and avoids providing initialized camera parameters. We analyzed the challenges of joint optimization, including coupling issue and intrinsics degenerated case. Based on this analysis, we design the framework with training sequence, calibration objects, and a calibration data acquisition strategy. Furthermore, we also provide a datasets with each image corresponding to different camera parameters, including both virtual and real-world scenes. The proposed method primarily tackles the substantial workload associated with acquiring multi-camera intrinsic and extrinsic parameters when employing multi-camera acquisition systems for 3D scene representation. Experiments demonstrate that proposed method achieves better rendering results and more accurate camera parameters. Code and models will be made available to the research community to facilitate reproducible research.
However, our method has several limitations: First, the multi-camera system requires a common field of view. Second, the images captured by these cameras must not contain image distortions. Third, in large-scale environments, $Pack1$ necessitates the use of large AprilTag cube to ensure the accuracy of parameter estimation. Based on our experience, in a multi-camera system covering an area of approximately 9.0m×$\times$6.0m×$\times$2.4m, using calibration blocks with a side length of 0.5m yields satisfactory reconstruction results. Employing smaller blocks significantly diminishes the precision and scale of the reconstructed scenes.
ACKNOWLEDGMENTS
The authors would like to thank Tianji Jiang, Xi Xu, Jiadong Tang, Zhaoxiang Liang, Xihan Wang, Dianyi Yang, bohan Ren, Yixian Wang, and all other members of the ININ Lab of the Beijing Institute of Technology for their contribution to this work.
@@ ACKNOWLEDGMENTS
Supplementary Materials
Supplementary performance in Real-World Datasets
In the original paper, EXP.4 in Sec 4.4 provided the rendering results for $Lab1$ Scene, while the rendering results for the other two scenes are shown in Fig. 13 and Fig. 14. The comparative methods and metrics setup are consistent with Fig. 12 in the original paper. From the rendering results of the $Lab2$ and $Lab3$ scenes, it can be seen that our method achieves competitive performance across different scenes, further demonstrating its effectiveness."
Supplementary Experiment A: Fix Step VS Global optimization
In this experiment, we aim to validate the effectiveness of the global optimization framework and compare its performance against the Fixed Step NeRF rendering (Intrinsic and extrinsic parameters estimation + NeRF, without other optimization). Fixed Step rendering simulates the typical steps involved in NeRF-based reconstruction using a multi-camera acquisition systems: Initially, camera intrinsic and extrinsic parameters are determined using calibration information. Subsequently, NeRF is trained using these fixed parameters to represent the 3D scene. Throughout the NeRF training phase, the camera parameters remain constant.
The experimental results are presented in Fig.15, Table 6, and Table 7. Regarding image rendering quality, in Fig.15, we randomly selected three scenes rendered in four different styles. The global optimization method stands out by producing sharper object boundaries and capturing finer object details. Table 6 shows the rendering performance for 4 styles, and Table 7 shows the camera parameters estimation results. Firstly, MC-NeRF shows a significant performance improvement in Array Style, where it achieves the best scores across all three evaluation metrics. Secondly, we observed that MC-NeRF consistently outperforms in all LPIPS metrics. However, when it comes to PSNR and SSIM, except for the Array scene, MC-NeRF often has lower scores. We explored the reasons, and the detailed analyses and experiments are discussed in SUPPLEMENTARY EXPERIMENT D.
Supplementary Experiment B: Stacking Techniques
In MC-NeRF, the NeRF component does not undergo significant modifications, the method for estimating camera parameters is derived from NeRF--- - and BARF. However, simply stacking methods cannot achieve the effect of MC-NeRF, mainly for two reasons: First, the parameter coupling issue, which has already been emphasized in the original manuscript. For this part, we provide additional AprilTag information to ensure the parameters can be decoupled. Second, the training process needs to follow the training sequence diagram proposed in the original manuscript (Fig. 6: Training Sequence Diagram). Here, we conducted comparative experiments by simply stacking the techniques of BARF and NeRF--- - and using the same camera parameter initialization process as MC-NeRF. The results are shown in Fig.16 and Table 8.
As shown in Table 8, Mix-Step represents the hybrid method of AprilTag + BARF + NeRF--- -, and the comparative results demonstrate that simply stacking techniques without adhering to the aforementioned conditions cannot achieve the desired results.
Supplementary Experiment C: NeRF with Off-the-Shelf Libraries
In addition to the comparison experiments with real-world datasets described in the original manuscript, we carried out further supplementary experiments to substantiate the performance of MC-NeRF in synthetic datasets. In these experiments, we employed camera parameters obtained using SFM (Structure from Motion) techniques from well-established libraries including COLMAP and Meshroom. These parameters are then used to train with NeRF. The count of usable images sourced from these libraries is presented in Table 9.
Table 9 demonstrates that with synthetic datasets, SFM-based methods exhibit unreliable performance, leading to a significant loss of image data. This can be attributed to the simplistic background textures prevalent in synthetic data, coupled with the inherently limited number of images available. Conversely, the estimation method using AprilTag does not encounter these issues, which is one of the reasons we opted for this method.
To ensure that the experimental results are meaningful, we restricted our training to scenes with more than 30 usable images. The rendering results illustrated in Fig. 17 and Table 10. From Fig. 17, it is seen that MC-NeRF achieves better rendering details, particularly in terms of object edge delineation. Furthermore, in quantitative results, MC-NeRF also outperforms the other two methods across all three metrics.
Supplementary Experiment D: PSNR and SSIM evaluation metrics analysis
Although MC-NeRF achieved better rendering details, it scored lower on both PSNR and SSIM evaluation metrics. We believe this is related to the alignment of object boundaries, and the reason behind this phenomenon is the absence of enforced ray alignment constraints for rendering in MC-NeRF. Specifically, the camera parameters of Fix-Step NeRF are constant and remain stable during the rendering stage, which maintains a fixed ray distribution, and all fitting computations contained within NeRF’s MLPs. It can be observed that to compensate for the ray distribution error resulting from inaccuracies in camera parameters, the rendered objects exhibit a multi-layer shadow structure. In Fig.18, the boundaries of the computer show noticeable shadows, while those of MC-NeRF appear much clearer and sharper. This multi-layer shadow structure has more advantages for pixel-to-pixel calculation methods like PSNR. If the clear-edged object generated by MC-NeRF is not aligned with the ground truth, the PSNR scores can be easily affected by the background, resulting in lower scores. Although MC-NeRF has some drawbacks in terms of boundary alignment, we prioritize high-quality rendering. Additionally, based on the rendering images, we believe that this alignment error is tolerable and unlikely to be readily noticeable.
Moreover, in this experiment, we noted that achieving higher-quality rendering results does not necessarily correlate with higher PSNR and SSIM scores. Because even when the estimation results for both intrinsic and extrinsic parameters are highly accurate, their values are often challenging to align perfectly with ground truth. This leads to slight shifts of objects in the rendered image, which have a notable impact when calculating pixel-wise differences for PSNR and SSIM. However, the LPIPS can still objectively reflect rendering performance even in such cases.
5.1   *Supplementary Experiment E: Unique-Camera VS Multi-Cameras
In this experiment, we aim to explore whether NeRF can simultaneously utilize multiple intrinsic parameters, a foundational prerequisite for our method. We select cameras with FOVs of 40, 50, 60, 70, and 80 degrees to generate a mixed dataset. NeRF is then trained on this dataset with the provided ground-truth intrinsic and extrinsic parameters. Simultaneously, we conduct training of NeRF with each FOV data independently to compare the rendering performance against NeRF trained on the mixed dataset. Hence, this experiment comprises a total of 6 datasets: one mixed dataset and five independent datasets, each containing 84 images with identical extrinsic parameters across all datasets.
The comparison results can be observed in Fig.19 and Table 11. The $Distance$ item in Table 11 indicates the absolute difference between their respective indicators. As evident from both Table 11 and Fig.19, the models trained with different datasets exhibit nearly equivalent performance. This observation has shown the capability of NeRF that can effectively process multi-camera images, when accurate camera intrinsic and extrinsic parameters are provided.
Figure 13: Rendering results of different methods in real-world dataset $Lab2$. In our comparative experiments, InstantSplat and CF Gaussian autonomously estimate camera parameters without any provided initializations. Methods such as CAMP, NeRFacto, and BARF, which are initialized using camera parameters provided by Colmap, are limited to optimizing only the extrinsic parameters. In contrast, SC-NeRF and NeRF--- -, also initialized with Colmap, are capable of optimizing both intrinsic and extrinsic parameters. Instant-NGP, which lacks camera parameter optimization capabilities, serves as a baseline reference. MC-NeRF utilizes auxiliary images for initialization, and it is also capable of optimizing both intrinsic and extrinsic parameters. The results demonstrate that our method achieved the competitive rendering performance among all compared methods.
Figure 14: Rendering results of different methods in real-world dataset $Lab3$. In our comparative experiments, InstantSplat and CF Gaussian autonomously estimate camera parameters without any provided initializations. Methods such as CAMP, NeRFacto, and BARF, which are initialized using camera parameters provided by Colmap, are limited to optimizing only the extrinsic parameters. In contrast, SC-NeRF and NeRF--- -, also initialized with Colmap, are capable of optimizing both intrinsic and extrinsic parameters. Instant-NGP, which lacks camera parameter optimization capabilities, serves as a baseline reference. MC-NeRF utilizes auxiliary images for initialization, and it is also capable of optimizing both intrinsic and extrinsic parameters. The results demonstrate that our method achieved the competitive rendering performance among all compared methods.
@@
TABLE 6: Quantitative rendering results of Fix-Step NeRF and MC-NeRF on all scenes. From the table, it is evident that while MC-NeRF does not perform well in terms of PSNR and SSIM, it significantly outperforms the Fix-Step method in LPIPS. We analyzed the reasons behind this unexpected result, which are discussed in detail in SUPPLEMENTARY EXPERIMENT D.
@@
TABLE 7: Camera parameters estimation results of Fix-Step NeRF and MC-NeRF on all scenes. $Loss\_K$, $Loss\_R$, and $Loss\_T$ represent the average loss of all cameras in the rendering scene for the intrinsic parameters, rotation vectors, and translation vectors, respectively.